4) Assume that IQ scores of college graduates are normally distributed. A researcher collects a random...

4)

Assume that IQ scores of college graduates are normally distributed. A researcher collects a random sample of 25 college graduates and tests their IQ. The sample had a mean IQ of 108 with a standard deviation of 15.6.

a) Find a 95% confidence interval for the true mean IQ for college graduates.

b) Provide the margin of error of the interval as your answer.

Round your answer to 1 decimal place.
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 108

sample standard deviation = s = 15.6

sample size = n = 25

Degrees of freedom = df = n - 1 = 25 - 1 = 24

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,24 = 2.064

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.064 * (15.6 / \sqrt 25)

= 6.4

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

108 - 6.4 < \mu < 108 + 6.4

101.6 < \mu < 114.4

(101.6 , 114.4)